2.12 g mixture of hexanoic acid and octanoic acid was dissolved in 100 ml of water and its 10 ml portion required 16.67 ml of a 0.1M standard alkali solution. determine the percentage composition of the mixture.

Grade:12

1 Answers

Sarthak gupta
3 years ago
Equivalents of Mixture is equal to equivalents of alkali medium
Let Hexanoic Acid be X gm and Octanoic Acid be 2.12 - X gm in mixture.
X / 116 * 1 = 2.12 - X /144 * 1 = 16.67 * 0.1 / 1000
X = 0.193
Hexanoic Acid -     9.1 %
Octanoic Acid -.      90.9 %